Solve the differential equation by separation of variables.

(x + 2) dx = (x + 3) sin y cos y dy
1
Expert's answer
2021-09-15T02:57:12-0400

"(x + 2) dx = (x + 3) \sin y. \cos y dy"

"\Rightarrow \frac{x+2}{x+3} dx=\frac{\sin2y}{2}dy"

Integrating both sides, we get

"\int\frac{x+2}{x+3} dx=\int\frac{\sin2y}{2}dy
\\\Rightarrow \int\frac{x+3-1}{x+3} dx=\frac{1}{2}\int\sin2ydy
\\\Rightarrow \int dx-\int \frac{1}{x+3}dx=-\frac{1}{2}.\frac{\cos 2y}{2}+c
\\\Rightarrow x-ln|x+3|=-\frac{\cos 2y}{4}+c"